OpenStudy (anonymous):

find the derivative of 10/x as x=-12, Please explain the steps!

OpenStudy (earthcitizen):

y= 10/x then find d(10x^-1)/dx

OpenStudy (anonymous):

How would I solve it? Would distribute? :0 @EarthCitizen

OpenStudy (earthcitizen):

recall the formula, that dy/dx = nx^(n-1)

OpenStudy (anonymous):

can you help use the formula? @EarthCitizen

OpenStudy (earthcitizen):

make 10/x be the inverse of x multiplied by 10 , which is 10x^(-1) then differentiate

OpenStudy (earthcitizen):

since y = 10x(-1) our formula says that dy/dx = \[d(10x^-1)/dx\] therefore \[dy/dx = nx ^{n-1}\]

OpenStudy (earthcitizen):

from our equation y = 10x^(-1) implies that n in the formula would be -1 therefore\[dy/dx = (-1)*10*x ^{-1-1}\]

OpenStudy (earthcitizen):

Hence \[dy/dx = -10x ^{-2}\] which can also be written as \[dy/dx ={-10}/{x^{2}}\]

OpenStudy (earthcitizen):

yh, check this site out it could help https://www.wolframalpha.com/input/?i=10%2Fx

OpenStudy (anonymous):

This is a question from a practice test I took that I got incorrect and they gave me a set of multiple choice answers and that's not one of the answer choices :0 @EarthCitizen

OpenStudy (earthcitizen):

yh, because you need to substitute x=-12 which should be that \[f^{\prime}(-12)= -10/{-12}^{2}\]

OpenStudy (earthcitizen):

Is -10/144 one of the answers in the options ?

OpenStudy (anonymous):

I had plugged it in and simplified and got 5/6 which is one of the answer choices but they said it was wrong

OpenStudy (anonymous):

no here are the options 72/5, 5/72, 6/5, 5/6

OpenStudy (anonymous):

btw, its -10/x at x=-12

OpenStudy (earthcitizen):

yh, it's should be 5/72 then ?

OpenStudy (anonymous):

how did you get that? XD

OpenStudy (earthcitizen):

yh, let's have a go at it again. \[y = -10x^{-1}\] therefore\[dy/dx = -1*-10*x^{-1-1} = 10x^{-2}\] at x = -12 implies that \[10x^{-2} = 10/x^{2} \implies f^{\prime}(-12) = 10/(-12)^2 = 10/144 = 5/72\]

OpenStudy (anonymous):

OHHHHHHHHHHHHH woooow!!! I'm sorry i see it now hahahaha

OpenStudy (earthcitizen):

yh, hehe

OpenStudy (anonymous):

Thank you :D

OpenStudy (earthcitizen):

no p, boss